40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 203 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  6875   Tue Jun 26 22:37:43 2012 yutaUpdateIOOenergized OMC stages

[Koji, Yuta]

We checked that PZTs between SRM and OMC (called OMC stage 1 and 2) is working.
Now we need them to be EPICS channels because they are not connected to digital world right now.

Background:
  For the IFO alignment, what we have been doing for last 2weeks is,

1. Align Y arm to Y end green and maximize green transmission
2. Use PZT2 to maximize TRY (PZT1 is not functioning well. PZT1 Y do a little, but X totally does nothing.)
3. Align BS and X arm to maximize TRX
4. Tune BS and ITMX so that reflection from both arms overlap at AS
5. Align X end green to that we can see bright(~250 uW) TEM00 at transmission

  However, we found that something (Y arm axis or Y end green?) has drifted horizontally and can't make Y green transmission and TRY high level at same time. Because PZT1 is not functioning well, it is hard to compensate beam translation.
  So, now what we have to do is to align Y arm to IR incident beam. That means, we either have to realign Y end green or forget about maximizing green transmission. I think I will leave green as it is for a while because calibration of the beatbox is going on and I want to proceed to PRC.
  Anyway, if we align IFO to the IR incident beam, we see clipping in the AS port. From the contrast measurement last night, we thought clipping comes from somewhere between BS and AS port. So, we need PZTs between BS and AS port working.

What we did:
  1. Turned on 24P 24N power supplies(Sorensen DCS33-33E) in AUX_OMC_SOUTH rack to supply power to AUX_OMC_NORTH rack. 18P 18N cables to OMC_NORTH was unplugged and used by the beatbox, so we reconnected them.

  2. Turned on KEPCO high voltage power supply to supply 150 V to the PZT driver, but it was not functioning well. So, we currently use Aligent HP 6209B instead. Its on the OMC_NORTH rack.

  3. PZT driver output to OMC stage 1 was unplugged. So, we plugged them.

  4. Opened PZT driver (LIGO-D060287), put some signal from Piezo_Drive_in(J4 in schematic), and checked beamspot at AS port is moving. The gain from Piezo_Drive_in to the output (hv_out) was ~20.

  5. We could avoid clipping by putting some offset to OMC stage 2 (or 1) in yaw. That means, the clipping comes from after OMC stage 2.

Conclusion:
  If we can control OMC stage 1 and 2, we can avoid clipping. So, we want them to be EPICS channels.

  6874   Tue Jun 26 01:30:13 2012 yutaSummaryGreen Lockingsimultaneous hold and release of the arm (aka two arm ALS)

To get the feeling of the master of IFO, I;

1. Stabilized both arm length using ALS.

2. Ran findIRresonance.py for both arms and find what offset gives me IR resonances.

3. Holded X arm to IR resonance, holded Y arm to IR resonance, and released both arms.

Below is the time series data of what I did.
ALSboth20120625_2withAS.png


Issues:
 - Currently ALS is not stable enough. It only stays for about few minutes. I think it is because of the bad alignment of green from each end.
 - We can't tell end green frequency is higher or PSL green frequency is higher. So, the sign of the servo filter sometimes flips.
 - Wobbliness of X end green transmission beam spot was from the ETMX oplev. When the oplev servo is on, it got more wobbly when X end table is opened. But when the oplev servo was off, wobbliness was same even if the presence of air flow.
 - MICH contrast in plot above seems like it somehow got better when two arms are at resonance by ALS. I think this is not real because reflection from both arms at AS port was not well aligned and beam was clipped. Koji and I measured contrast of FPMI and MI(ETMs misalined), and they were 99.6 % and 99.9 % respectively. Beam clipping seems like it comes from some where between BS to AS port. We need to figure out where and fix this.

Things need to be done to make ALS more concrete:
 - Align Y end green beam!
 - Look into Y end green frequency servo
 - How do we hand-off servo using ALS to IR lock?
 - Noise budgeting for new phase tracker scheme
 - Linearity check of the beat box and phase tracker

  6873   Tue Jun 26 00:52:18 2012 yutaUpdateComputer Scripts / ProgramsPMC locker

Quote:

I thought we rewrite auto lockers once per year, but this time it took us only a month. I wrote it for PMC on May 24. Is it not working?

I know.
I just wanted to use pyNDS for this kind of scanning & locking situation.
c1ioo was down for the weekend and I couldn't test my script for ALS, so I used it for PMC.

But I think PMClocker.py can relock PMC faster because it can sweep C1:PSL-PMC_RAMP continuously and can get continuous data of C1:PSL-PMC_PMCTRANSPD.

  6872   Mon Jun 25 21:54:52 2012 DenUpdateComputer Scripts / ProgramsPMC locker

Quote:

I made a python script for relocking PMC.
It currently lives in /opt/rtcds/caltech/c1/scripts/PSL/PMC/PMClocker.py.

I thought we rewrite auto lockers once per year, but this time it took us only a month. I wrote it for PMC on May 24. Is it not working?

Could someone make it more clear why some scripts are written on bash, others on sh or python? I think we should elaborate a strict order. Masha and I can work on it if anyone else considers this issue as a problem.

  6871   Mon Jun 25 17:48:27 2012 yutaUpdateComputer Scripts / Programsscript for finding IR resonance using ALS

I made a python script for finding IR resonance using ALS. It currently lives in /opt/rtcds/caltech/c1/scripts/ALS/findIRresonance.py.

The basic algorism is as follows.

1. Scan the arm by putting an offset to the phase output of the phase tracker(Step C1:ALS-BEAT(X|Y)_FINE_OFFSET_OFFSET by 10 deg with 3 sec ramp time).

2. Fetch TR(X|Y) and OFFSET online data using pyNDS during the step.

3. If it finds a tall peak, sets OFFSET to the value where the tall peak was found.

4. If tall peak wasn't found, go back to 1 and step OFFSET again.

The time series data of how he did is plotted below.
I ran the script for Y arm, but it is compatible for both X and Y arm.

findIRresonance20120625.png

  6870   Mon Jun 25 16:21:10 2012 KojiSummaryIOOSelection of motorized mirror mounts

I am considering to have 3 to 6 motorized optical mounts at the PSL and end tables for remote beam steering.

Question 1:

Was there any issue on the PI 3-axis PZT on the PSL?
Why was it disabled (even before the PSL upgrade)?


Question 2:

Do we need two mount at a place? Or we do have one instead?

- Comparing the distance of the steering mirrors and that from the steering mirror to the cavity waist, induced shift
is mostly cancelled by angle adjustemnts of a either of the mounts.
i.e. Induced misalignments by the steering mirrors are nearly degenerated.

We need to move two steering mirrors only for the initial installation, but any drift felt by a cavity can be compensated by a single mirror.

Question 3:
Do we like PI-style 2 or 3 axis PZT mount with analog inputs on the HV amp?
Or do we like "Newport Agilis" style controller with USB connection?

Any opinion?

  6869   Mon Jun 25 15:19:07 2012 YaakovUpdatePEMAdded microphone channels, moved accelorometer channels

Jenne and I renamed the mic channels Den created (elog 6664) to MIC_1, MIC_2, etc from the original accelerometer names to keep things clear. We then added 6 new channels (22-27) for the accelerometers, named ACC_MC1_X, Y, Z, ACC_MC2_X, Y, Z, etc. (See the screenshot below). We also added a DAQ channel block and listed out the IN1 channel of all the sensors. We compiled and started the model, and checked that all the channels were there in DataViewer.

channels.png

  6868   Mon Jun 25 15:07:49 2012 yutaUpdateIOOMC beam spot trend

I adjusted MC WFS offsets using /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets.
Beam spot positions on MC mirrors came back to where it was past few weeks. See the trend below. Trend sometimes shows huge jump, but it's just a bad measurement caused by unlock of MC during the measurement.

I ran /opt/rtcds/caltech/c1/scripts/ASS/MC/mcassMCdecenter to measure beam spot whenever I feel like it (see elog #6727).
Beam spot doesn't move so much (~0.2 mm in standard deviation), which means incident beam from PSL table is quite stable.


MCdecenter.png

  6867   Mon Jun 25 11:27:13 2012 JamieSummaryComputersc1ioo is down

Quote:

Quote:

I tried to restart c1ioo becuase I can't live without him.

I couldn't ssh or ping c1ioo, so I did hardware reboot.
c1ioo came back, but now ADC/DAC stats are all red.

c1ioo was OK until 3am when I left the control room last night. I don't know what happened, but StripTool from zita tells me that MC lock went off at around 4pm.

 c1ioo was still all red on the CDS status screen, so I tried a couple of things.

mxstreamrestart (which aliases on the front ends to sudo /etc/init.d/mx_stream restart) didn't help

sudo shutdown -r now didn't change anything either....c1ioo came back with red everywhere and 0x2bad on the IOP

eventually doing as Jamie did for c1sus in elog 6742, rtcds stop all, then rtcds start all fixed everything.  Interestingly, when I tried rtcds start iop, I got the error
Cannot start/stop model 'iop' on host c1ioo, so I just tried rtcds start all, and that worked fine....started with c1x03, then c1ioo, then c1gcv.

There seems to be a problem with how models are coming up on boot since the upgrade.  I think the IOP isn't coming up correctly for some reason, which is then preventing the rest of the models from starting since they depend on the IOP.

The simple way to fix this is to run the following:

ssh c1ioo
rtcds restart all

The "restart" command does the smart thing, by stopping all the models in the correct order (IOP last) and then restarting them also in the correct order (IOP first).

  6866   Mon Jun 25 11:23:14 2012 JenneUpdateComputerstdsavg not working

Quote:

LSCoffsets script, and any others depending on tdsavg will not work until this is fixed.

 LSCoffsets is working again. 

tdsavg (now, but didn't used to) needs "LIGONDSIP=fb" to be specified.  Jamie just put this in the global environment, so tdsavg should just work like normal again.

Also, the rest of the LSCoffsets script (really the subcommand offset2) was tsch syntax, so I created offset3 which is bash syntax.

Now we can use LSCoffsets again.

  6865   Mon Jun 25 10:35:59 2012 JenneSummaryComputersc1ioo is down

Quote:

I tried to restart c1ioo becuase I can't live without him.

I couldn't ssh or ping c1ioo, so I did hardware reboot.
c1ioo came back, but now ADC/DAC stats are all red.

c1ioo was OK until 3am when I left the control room last night. I don't know what happened, but StripTool from zita tells me that MC lock went off at around 4pm.

 c1ioo was still all red on the CDS status screen, so I tried a couple of things.

mxstreamrestart (which aliases on the front ends to sudo /etc/init.d/mx_stream restart) didn't help

sudo shutdown -r now didn't change anything either....c1ioo came back with red everywhere and 0x2bad on the IOP

eventually doing as Jamie did for c1sus in elog 6742, rtcds stop all, then rtcds start all fixed everything.  Interestingly, when I tried rtcds start iop, I got the error
Cannot start/stop model 'iop' on host c1ioo, so I just tried rtcds start all, and that worked fine....started with c1x03, then c1ioo, then c1gcv.

  6864   Mon Jun 25 08:21:40 2012 steveUpdateGeneralAC power disturbance on Sat

Quote:

I was notified by CIT Utilities that there was a power surge or short power outage this after noon.

Lab conditions are normal:  c1ioo is down.  The south arm AC was off......I turned it back on.

 

             CALIFORNIA INSTITUTE OF TECHNOLOGY

                 FACILITIES MANAGEMENT

                 UTILITY & SERVICE INTERRUPTION

 

**PLEASE POST**

 

Building:         CAMPUS

 

Date:             Saturday, June 23, 2012

 

Time:             3:46 P.M.

 

Interruption:     Electrical Power Disturbance

 

Contact:          Tom Brennan, x-625-395-4984    

 

*The City of Pasadena Water & Power Department had a 34,000-volt line event on Saturday June 23 at 3:46 p.m.  This caused a city wide disturbance on the power grid.  The Campus did not lose electrical power.  However, the disturbance may have affected sensitive electronic equipment.

(If there is a problem with this Interruption, please notify the Service Center X-4717 or the above Contact as soon as possible.

If no response is received we will proceed with the interruption.)

         

                        Jerry Thompson,

                        Interim Director of Campus Operations & Maintenance

 

Attachment 1: powerdisturbed.png
powerdisturbed.png
  6863   Sun Jun 24 23:42:31 2012 yutaUpdateComputer Scripts / ProgramsPMC locker

I made a python script for relocking PMC.
It currently lives in /opt/rtcds/caltech/c1/scripts/PSL/PMC/PMClocker.py.

I think the hardest part for this kind of locker is the scan speed. I could make the scan relatively fast by using pyNDS.
The basic algorithm is as follows.

1. Turns off the servo by C1:PSL-PMC_SW1.

2. Scans C1:PSL-PMC_RAMP using ezcastep.bin. Default settings for ezcastep is

ezcastep.bin C1:PSL-PMC_RAMP -s 0.1 0.01 10000

So, it steps by 0.01 for 10000 times with interval of 0.1 sec.

3. Get C1:PSL-PMC_PMCTRANSPD and C1:PSL-PMC_RAMP online 1 sec data using pyNDS.

4. If it finds a tall peak in C1:PSL-PMC_PMCTRANSPD, kills ezcastep.bin process, sets C1:PSL-PMC_RAMP to the value where the tall peak was found, and then turns on the servo.

5. If tall peak wasn't found, go back to 3 and get data again.

6. If C1:PSL-PMC_RAMP reaches near -7 V or 0 V, it kills previous ezcastep.bin process and turns the sign of the scan.

I tested this script several times. It sometimes passes over TEM00 (because of the dead time in online pyNDS?), but it locks PMC with in ~10 sec.
Currently, you have to run this to relock PMC because I don't know how to make this an autolocker.

I think use of pyNDS can be applied for finding IR resonance using ALS, too.
I haven't checked it yet becuase c1ioo is down, but ALS version lives in /users/yuta/scripts/findIRresonance.py. ALS may be easier in that we can use fast channels and nice filter modules.

Other scripts:
 I updated /opt/rtcds/caltech/c1/scripts/general/toggler.py. It now has "lazymode". When lazymode, it toggles automatically with interval of 1 sec until you Ctrl-c.

 Also, I moved damprestore.py from my users directory to /opt/rtcds/caltech/c1/scripts/SUS/damprestore.py. It restores suspension damping of a specified mirror when watchdog shuts down the damping.

  6862   Sun Jun 24 00:10:45 2012 yutaUpdateGreen Lockingcurrent beat electronics

I moved amplifiers for beat PD at PSL table to 1X2 rack. Current beat setup from PD to ADC is shown below. Setup for X beat and Y beat are almost the same except for minor difference like cable kind for the delay line.

Currently, DC power for amplifiers ZHL-1000LN+ is supplied by Aligent E3620A.
I tried to use power supply from the side of 1X1 rack, but fuse plug(Phoenix Contact ST-SI-UK-4) showed red LED, so I couldn't use it.
Measured amplification was +25 dB for 10-100 MHz.

Measured gain from RF input to monitor output of the beat box was ~ -1 db for 10-100 MHz.

beatsetup20120623.png

  6861   Sat Jun 23 19:57:22 2012 yutaSummaryComputersc1ioo is down

I tried to restart c1ioo becuase I can't live without him.

I couldn't ssh or ping c1ioo, so I did hardware reboot.
c1ioo came back, but now ADC/DAC stats are all red.

c1ioo was OK until 3am when I left the control room last night. I don't know what happened, but StripTool from zita tells me that MC lock went off at around 4pm.

  6860   Sat Jun 23 18:44:15 2012 steveUpdateGeneralpower surge has no effect on the lab

I was notified by CIT Utilities that there was a power surge or short power outage this after noon.

Lab conditions are normal:  c1ioo is down.  The south arm AC was off......I turned it back on.

  6859   Sat Jun 23 02:29:18 2012 yutaUpdateGreen LockingX arm mode scan results

X arm finesse is 416 +/- 6, mode-matching ratio is 91.2 +/- 0.3%

I did mode scan for X arm just like we did for Y arm (see elog #6832)

Servo design:
  Servo filters are as same as Y arm.
  UGF and phase margin of X arm ALS are 100 Hz and 14 deg.
  For phase tracking loop, they are 1.5 kHz and 56 deg.

Raw data from the mode scan:
XarmScan20120623.png


Fitted peaks and finesse:

fine8FSRscanXarm.png

By taking the average,

F = 416 +/- 6 (error in 1 sigma)
(For Y arm, it was 421 +/- 6. See elog #6832)


Mode matching ratio:
 From X arm 8FSR measurement using phase tracker, peak heights are

TEM00 0.834, 0.851, 0.854, 0.852, 0.876, 0.850, 0.855, 0.878
TEM01 0.031, 0.031, 0.017, 0.017, 0.009, 0.014, 0.009, 0.011
TEM02 0.053, 0.052, 0.057, 0.058, 0.061, 0.060, 0.061, 0.059
TEM03 0.011, 0.010, 0.010, 0.007, 0.006, 0.005, 0.006, 0.005

 So, the mode-matching ratio is

MMR = 89.7%, 90.1%, 91.0%, 91.2%, 92.0%, 91.4%, 91.8%, 92.1%

 By taking the average,

MMR =  91.2 +/- 0.3 (error in 1 sigma)
(for Y arm, it was 86.7 +/- 0.3 %. See elog #6828)


Discussion:
 - Mode matching ratio for both X and Y arm is ~90%, which is not so great, but OK. It seems like there's no huge clipping or mode-mismatch from MC to ITMs. I think we should go next for PRMI investigation.

 - Measured finesse seems too low compared with the design value 450. If we believe power transmission of ITM and ETM are 0.0138 and 1.37e-5, the measured finesse tells you that there's ~0.1% loss(F = 2*pi/(T_{ITM}+T_{ETM}+T_{loss})). We need some evaluation for the linearity of the sweep, before concluding that there's 0.1% loss for each arm. Using FINE_I/Q signal for calibration, or installing frequency divider for monitoring actual beat frequency would help.


Things to do for the beat setup:

 - Amplifiers after beat PDs shouldn't be on the PSL table. Move them near the beatbox.
 - Install DC PD (and camera?) at un-used port of the beat BS for monitoring green transmission power.
 - Make nice MEDM screens for our new phase tracking ALS.
 - Make a script to sweep arm length with ALS and find IR resonance.
 - Look into X end table. Beam spot of the X green transmission is wiggly when X end table is opened and there's air flow.

  6858   Fri Jun 22 20:58:15 2012 JenneUpdateGreen LockingCalibrated POX spectra - Xarm stabilized by ALS

[Yuta, Jenne, Koji]

We stabilized the Xarm using the ALS and took a spectrum of POX as our out of loop sensor.  We used the calibration from elog 6841 to go from counts to meters.

We find (see attached pdf) that the RMS is around 60pm, dominated by 1Hz motion.

 

In other, related, news, I took out the beam pipe connecting the AP and PSL tables and covered the holes with foil.  This makes it much easier and faster to get to the X beat setup for alignment.  Eventually we'll have to put it back, but while the AUX laser on the AP table is not being used for beating against the PSL it'll be nice to have it out of the way.

Attachment 1: POX11_I_ERR_calib_residualCavityMotion_better.pdf
POX11_I_ERR_calib_residualCavityMotion_better.pdf
  6857   Fri Jun 22 20:00:14 2012 JamieOmnistructureElectronicstwo RG-405 cables ran from 1X2 rack to control room

[Yaakov, Eric, Jenne, Yuta]

Two of our surfs, Yaakov and Eric, pulled two unused RG-405/SMA cables that had been running from 1X2 to (mysteriously) 1Y2 racks.  They left the 1X2 end where it was and pulled the 1Y2 end and rerouted it to the control room.  We labeled both ends appropriately.

The end at 1X2 is now plugged into a splitter that is combining the RF input monitor outputs for the X and Y beatbox channels, so that we can watch the beat signals with the HP8591 in the control room.

  6856   Fri Jun 22 19:52:47 2012 JamieUpdateComputersottavia reconfigured as CDS workstation

ottavia has been reinstalled with Ubuntu 10.04 LTS, and has been configured as a CDS workstation.

I have been maintaining a script that takes a stock 10.04 install and configures it as a workstation.  I've attached it here, but it lives at:

/users/controls/workstation-setup.sh

The script is designed to be idempotent, i.e. it can be run on a machine that has already been configured and it will either have no affect or update.

Attachment 1: workstation-setup.sh
#!/bin/bash

# This is a CDS workstation setup script for Ubuntu 10.04 It is
# idempotent, so running it on an already configured workstation is
# fine.

########################################
### controls

if [[ $(id -u controls) != 1001 ]] ; then
... 111 more lines ...
  6855   Fri Jun 22 17:51:04 2012 JenneUpdateCamerasGreen Trans camera repositioning attempt

[Yuta, Jenne]

We tried to find a different place, not in the main green transmitted beam path, to place the trans camera for the green beams.  There is a little bit of leakage through the 3 high reflector mirrors which steer the beams from the direction when they first come out of the chamber over to the main green beat setup.  2 of these mirrors have virtually no space behind them for a camera (the first one the green beams encounters is right next to the EOM mount, and the 2nd one is pretty close to the Input Pointing QPDs.  We can potentially use the beam leaking through the 3rd steering mirror, if the camera is very close to the edge of the table (so that the camera isn't blocking the IR input pointing beams), but the X beam is so dim as to be nearly impossible to see, even when TEM00.  This precludes the point of the camera, which is to see the modes when we're aligning the beams.  (X power on the PSL table is pretty high - 330uW measured today, but those mirrors must transmit the Y beam's polarization more than the X beam's.)

Our other thought was to use one of the secondary beams coming out of the chambers.  This is kind of Mickey Mouse, but we thought that since this is just a camera to see the modes, as opposed to a PD, maybe it's okay.  This is a moot point however, since the secondary and tertiary beams (due to the wedge of the window) are clipped for the Y green.  We closed the PSL shutter then removed the beam pipe between the PSL table and the chamber so I could look inside. 

It looks to me like the main green transmitted beams are exiting through the window several inches from any edge, so they're definitely not clipping.  But the reflection from the window back into the chamber is hitting some optic.  The X green is hitting the face of the optic, while the Y green is hitting the edge of the optic and part of the mount.  The reflections from this mount then go back toward the chamber window and out toward the PSL table.  This isn't a big deal for the camera situation - we'll just use the leakage from one of the steering mirrors somehow, but it does mean that there is some green light reflected back onto an IR mirror, and potentially causing grief.  I didn't look to see if the mirror it's hitting is the 1st in-vac IR steering mirror (I don't think so) or something in the OMC / AS path (I think it's something here), but either way, we could be making trouble for ourselves.  We should try to dump the reflection from the window when we vent.  Jamie has put it on the List.

We replaced the beam pipe between the PSL table and the chamber before opening the shutter on the laser.  We are currently sticking with the plan of putting the camera in the main green trans path for initial alignment, then removing it for the rest of the work.

  6854   Fri Jun 22 13:37:17 2012 JenneUpdateComputersfb lost connection

...Perhaps related to the fact that Jamie is copying a lot of stuff over the network to back up Ottavia before converting her to Ubuntu, perhaps totally independent. 

After restarting the daqd, c1lsc was the only computer whose mx_stream came up on its own.  I restarted c1sus. c1ioo, c1iscey, c1iscex by hand.

  6853   Fri Jun 22 10:52:18 2012 yutaUpdateGreen Locking2 arm ALS - Success!!!!

Answers to questions from Koji.

Are these correct?

1. It is a nice work.

Correct, of course!

2. This is not locking, but stabilization of the both arms by ALS.

Correct.

3. We now have the phase trackers for both arms.

Correct.

4. There is no coarse (i.e. short) delay line any more.

Correct. No coarse, only fine delay line (30m) with the phase tracker.

5. The splitters after the PDs are reducing the RF power to Beat-box.
Actually there are RF monitors on Beat-box for this purpose, but you did not notice them.

Oh, yes. But distance between beatbox and spectrum analyzer in the control room is longer than distance between BBPD on PSL table and the spectrum analyzer. We were too lazy to do cabling, but maybe we should.

6. c1ioo channel list 
https://wiki-40m.ligo.caltech.edu/CDS/C1IOO%20channel%20list
has to be updated.

Yes, we will.

7. Video can be uploaded to Youtube as Mike did at http://nodus.ligo.caltech.edu:8080/40m/6513

We didn't, but we can.

  6852   Fri Jun 22 03:37:42 2012 KojiUpdateGreen Locking2 arm ALS - Success!!!!

Are these correct?

1. It is a nice work.

2. This is not locking, but stabilization of the both arms by ALS.

3. We now have the phase trackers for both arms.

4. There is no coarse (i.e. short) delay line any more.

5. The splitters after the PDs are reducing the RF power to Beat-box.
Actually there are RF monitors on Beat-box for this purpose, but you did not notice them.

6. c1ioo channel list
https://wiki-40m.ligo.caltech.edu/CDS/C1IOO%20channel%20list
has to be updated.

7. Video can be uploaded to Youtube as Mike did at http://nodus.ligo.caltech.edu:8080/40m/6513

  6851   Fri Jun 22 02:21:57 2012 JenneUpdateGreen Locking2 arm ALS - Success!!!!

[Yuta, Jenne]

We locked both arms using the ALS system simultaneously!  Hooray!

Video of spectrum analyzer during lock acquisition of both beats is attached.

Jamie is super awesome, since he fixed us up a beatbox speedy-quick.  Thanks Jamie!!  speedy_gonzales-5257.jpg

 Details:

1:  Aligned PSL green optics

     1.1:  We added an amplifier of ~20dB after the X beat PD (more Xgreen power on the PSL table so the signal was ~3dB higher than Y, so required less amplification).  The ~24dB amplifier is still in place after the Y beat PD.  Both beat signals go to a splitter after their amplifiers.  One side of each splitter goes to one of the channels on the beatbox.  The other side of each splitter goes to a 3rd splitter, which we're using backwards to combine the 2 signals so we can see both peaks on the spectrum analyzer at the same time.

2:  Found both beat notes

     2.1:  Y beat was easy since we knew the temps that have been working for the past several days

     2.2:  X beat was more tricky - the last time it was locked was the end of February (elog 6342)

         2.2.1:  We found it by adjusting the PSL laser temp nearly the full range - DC Adjust slider was at 8.8V or so (Y beat was found with the slider at ~1.1V tonight)

          2.2.2:  We then walked the beat around to get the PSL temp back to "normal" by moving the PSL temp, then compensating with the Xend laser temp, keeping the beatnote within the range of the spectrum analyzer.

          2.2.3:  Fine tuned the temps of all 3 lasers until we had 2 peaks on the analyzer at the same time!!

               2.2.3.1:  Yend - measured Temp=34.14 C, thermal Out of Slow servo=29820

               2.2.3.2:  Xend - displayed temp=39.33 C, thermal Out of Slow servo=5070

               2.2.3.3: PSL - displayed temp=31.49 C, Slow actuator Adjust=1.100V

3:  Locked both arms using ALS!!

     3.1:  We were a little concerned that the Xarm wasn't locking.  We tried switching the cables on the beatbox so that we used the old channels for the Xarm, since the old channels had been working for Y.  Eventually we discovered that the input of the filter module for ETMX's POS-ALS input was OFF, so we weren't really sending any signals to ETMX.  We reverted the cabling to how it was this evening when Jamie reinstalled the beatbox.

          3.1.1:  We need to sort out our SUS screens - Not all buttons in medm-land link to the same versions of the SUS screens!  It looks like the ALS screen was modified to point the ETMY button to a custom ETMY SUS screen which has the ALS path in the POS screen, along with LSC and SUSPOS.  There is no such screen (that I have found) for ETMX.  The regular IFO_ALIGN screen points to the generic SUS screens for both ETMY and ETMX, so we didn't know until Yuta searched around for the filter bank that the ALS input for ETMX was off.  We just need to make sure that all of the screens reflect what's going on in the models.

     3.2:  See the video attached - it shows the beat peaks during locking!!! (how do I embed it? right now you have to download it)

          3.2.1:  First you will see both peaks moving around freely

          3.2.2:  Then X arm is locked briefly, then unlocked

          3.2.3: Y arm is locked, steadily increasing gain

          3.2.4:  X arm is locked, so both arms locked simultaneously

          3.2.5:  Yuta clicked a button, accidentally unlocking the Xarm

4:  The transmission of the X arm was not so great, and both of our green beams (although X green especially) were no longer nicely aligned with the cavities.  Yuta tried to align the X arm to the X green, but it's bad enough that we really need to start over with the whole IFO alignment - we leave this until tomorrow.  Since we didn't have any good IR transmission, we didn't bother to try to find and hold the Xarm on IR resonance using ALS, so we didn't measure a POX out of loop residual cavity motion spectrum.  Again, tomorrow. 

Attachment 1: P6210140.AVI
  6850   Thu Jun 21 20:07:18 2012 JamieUpdateGreen LockingImproved beatbox returns

I've reinstalled the beatbox in the 1X2 rack.  This improved version has the X and Y arm channels stuffed, but just one of the DFD channels (fine) each.

I hooked up the beat PD signals for X and Y to the RF inputs, and used the following two delay lines:

  • X: 140' light-colored cable on spool
  • Y: 30m black cable

The following channel --> c1ioo ADC --> c1gcv model connections were made:

  • X I  --> SR560 whitening --> ADC 22 -->  X fine I
  • X Q --> SR560 whitening --> ADC 23 --> X fine Q
  • Y I --> SR560 whitening --> ADC 24 --> Y fine I
  • Y Q --> SR560 whitening --> ADC 25 --> Y fine Q

The connections to the course inputs on the ALS block were grounded.  I then recompiled, reinstalled, and restarted c1gcv.  Functioning fine so far.

 

  6849   Thu Jun 21 15:36:51 2012 yutaUpdateLockingX arm alignment

I aligned X arm so that the beam spot comes roughly on the center.

1. Use ITMX and ETMX (mainly ITMX) to make beam spot come on center of the optic using eyeball.

2. Use ETMX and BS to maximize TRX power (reached ~ 0.85)

3. Aligned green optics on X end. Transmission of X green measured at PSL table is now 255 uW and TEM00 has the most power.

It was not easy to increase X green transmission more because beam spot on green transmission PD is wiggly when X end table is opened. When closed, wiggliness is about the same for Y green and X green.
Turning off HEPA on the X end didin't helped, but there must be something bad in the X end table. If we couldn't figure out why, let's wait for PZTs to come for end tables.

Considering the laser power is different(X end 1 W, Y end 700 mW), X green transmission should reach ~400 uW. But I think we should go on to X beat search.

I placed green shutter for X end back for convenience. I put some spacers to adjust its height and avoid beam clipping.


[Steve, Yuta]

What causing wiggly X green transmission was the air flow from the air conditioner. When we turned it off, beam spot motion became quiet. Air flow from HEPA was not effecting much.

  6848   Thu Jun 21 15:00:55 2012 steveUpdateSAFETYSURFs 2012 safety training

Masha, Eric, Yaakov, Liz and Sasha received 40m specific basic safety training.

Attachment 1: surfs2012.JPG
surfs2012.JPG
  6847   Thu Jun 21 12:56:49 2012 yutaUpdateLockingETMX 1064 trans camera

Quote:

[Jenne, Yuta]

We made ETMXT camera working.

 Xarm_EndTableLayout_NewTransCamera.png

Here's the new end table layout, for the transmitted IR stuff.

  6846   Thu Jun 21 12:13:35 2012 JenneUpdatePEMflow bench must be running all times

Quote:

The south end flow bench HEPA filter should be run all times. You can turn it off for a measurement or two but remember we are storing clean optics there.

The zero count bench will reach  room particle count  ~ 10,000 in one minute.

 My bad.  I turned it off last night to see if it would help make the Xgreen more stable, and then when I woke up this morning I realized that I had forgotten to turn it back on.  Bad Jenne.

  6845   Thu Jun 21 09:10:08 2012 steveUpdatePEMflow bench must be running all times

The south end flow bench HEPA filter should be run all times. You can turn it off for a measurement or two but remember we are storing clean optics there.

The zero count bench will reach  room particle count  ~ 10,000 in one minute.

  6844   Thu Jun 21 09:01:18 2012 steveUpdateSTACISthe resurrection of STACIS -surf

There is a small wood cabinet under the south end flow bench, labeled STACIS.

Unit is complete with extension cards and cables.

Attachment 1: IMG_1364.JPG
IMG_1364.JPG
  6843   Thu Jun 21 02:57:16 2012 JenneBureaucracyGeneral40m Action Items

PRIORITY ITEMS:

ASS doesn't run on Ubuntu!
Put beatbox back, simultaneous arm ALS
Input matricies, output filters to tune SUS.  check after upgrade.
POX11 whitening is not toggling the analog whitening???
Look into PMC PZT drift - PZT failing? Real MC length change?
Vent planning / organization

THE FULL LIST:

General

     Audio system for the signals!!!! Even a crappy one!

SUS

     Input matricies, output filters to tune SUS.  Check after upgrade.

IFO

     Fix occasional common-mode power transient in the arm transmissions. Probably an alignment thing. Would ISS help?

     Drift of the green incident axis -> Assess the amount of the drift / replace the mount

     Calibration of POP22 / AS110

     PMC/IMC/ARM characterization (loss, finesse, reflectivity, etc)

     Arm cavity sweeps, mode scan

     Align AS OSA (others?)

     Investigate PRMI glitches, instability

     PZT or Picomotor mounts for PSL/ALS beams

     ALS on the both arm simultaneously / common / diff ALS scripts

     Measure green locking (Aux laser to arm) transfer functions, residual spectra

     Measure oplev spectra while locking Xgreen - see if the optics are particularly noisy

     Measure Xarm residual motion using POX while ALS is engaged.

     Fix Vio2 filter modules on SUS 

     Switch power supply for amplifiers of beatnote signal to rack power

     Add temp sensors for end lasers to CDS slow channels

     Put windows / pickoffs on PSL table for (a) green trans camera, (b) GTRY, (c) GTRX

CDS

     Capture OSA signals in CDS (the 'scope TDS1001B has a USB port in the back for connecting to the computer)

     Transmon (arms) for high and low power

     POX11 whitening is not toggling the analog whitening???

     Install guardians to monitor EPICS values

Electronics

     Actuator noise level characterization (coil driver response in V/m & coil driver noise level V/rtHz)

     Improvement of POP22/110/AS110 RF circuits?

MEDM

     Complete 40m overview screen - everything should be clickable with pseudo 3D icons

     Script to generate a MEDM tree

     Resurrect MEDM snapshots

     New ! buttons on every screen, include wiki page

     Add all screens to svn

SCRIPT

     Daily diagnosis of the MC spot positions (there must be something already...)

     Daily/occasional adjustment of the incident axis on the MC

     Panic button on Watchdog screen isn't working on Ubuntu

     OPLEV/OSEM trending script before the IFO work for diagnosis. Put into 40m summary screen.

     Auto-locker for arms

     Auto-locker for PSL things

     Diagnostic script for CDS - mx_stream, other stuff.

     Make sure scripts are all svn-ed

Video

     If each video screen has a caption, that would be great

     GUI interface of "videoswitch"

Ubuntu vs. CentOS

     Upgrade Ottavia to Ubuntu, make sure connect to DTT, Dataviewer, AWG.

MM

     IPPOS beam measurement

     AS beam measurement (if beam is bright enough)

     Mode matching calculations, sensitivity to MC waist measurement errors, PRM position

     Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT

     Use sensoray to capture, measure beam mode at AS, POP

Stray Light

     Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras

     Align AUX laser into dark port

     Assemble in-vac beam dumps - how many do we need?

OAF

     OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive

     Static-only OAF noise budget (Adaptive noise budget as next step)

     Script for daily / weekly re-calculation of Wiener, post to elog if need changing

Tip Tilts

     Prepare electronics for TTs (coil drivers)

      In-air TT testing to confirm we can control / move TTs before we vent

     Connect TTs to digital system and controls, lay cables if needed

     Determine whether we need to add a new flange to OMC chamber

RF Photodiodes

     Opto Energy diode laser - purchase

     Set everything up

     Demod board for AS110 - so we can also have POP110?

  6842   Thu Jun 21 01:58:29 2012 JenneUpdateGreen LockingXgreen preparations

[Yuta, Koji, Jenne]

Lots of small things happened tonight, in preparation for having both arms' ALS working simultaneously.

1. Xarm aligned in IR

     1.1 ETMX oplev centered

2. Xgreen coarsely aligned to Xarm

3. X beat setup on PSL table resurrected. 

     3.1 Steering optics for both X and Y green (before PBS) were touched to fix clipping Xgreen on some of the first mirrors after the light exits the chambers.

     3.2 Xgreen aligned to beat PD

     3.3 PSL green waveplate rotated so ~half of the light goes to X beat, other ~half goes to Y beat (recall we had rotated the polarization so we had max light on the Y beat PD a few weeks ago).

          3.3.1 Now we have ~80uW of PSL green going to each beat PD.

     3.4 PSL green aligned to X beat PD

          3.4.1 Replaced mount for mirror between PBS (which splits PSL green light) and BS (which combines PSL green and X green) so that I could get the alignment correct without having to use the full range of the knobs on the mount.

     3.5 Realigned (coarsely) Ygreen to Y beat PD - the mirrors just after the chambers had been touched, so Y green was no longer directly on the PD.  This will need to be done more finely when we're ready to lock the Yarm again.

     3.6 Dedicated cables for the DC of each beat PD were put in place, so we have those in addition to the DC transmission PDs which we are putting in temporarily each time we align the green to the cavities.  Some mystery unused cables that were running under the PSL table were removed.  The power for the X beat PD was rerouted so that it's much closer to the actual diode, and out of the way. 

4. Better alignment of X green to X arm.

     4.1 Put Green Transmission camera into place

     4.2 Noticed that the X green spot on the transmission camera is not nearly as steady as the Y green.  Increased the gain of the X green refl PD on the end table to see if it helped the spot be more steady, but it's still very wiggly.  We reverted the gain to what it was.  We need to fix this!!!!

    4.3 Removed camera, looked at X transmission DC (PD is temporarily in front of the beat PD), tried to increase the transmission.

     4.4 Aligning the green to the X arm has been really tough - there were a few more iterations of camera then DC PD.

     4.5 Measured X green power on the PSL table - 02 mode was ~150uW.  The 00 mode is still not very stable, which is frustrating, although we have a reasonable amount of power transmitted.   

     4.6 The X end green shutter was moved out of the beam path since the green beam was clipping while going through the shutter.  We need to put it back now that the beam is pretty much aligned.  The beam size and the aperture are roughly the same, so we should look to see if there is a different place on the table where the beam is a little smaller, where we can put the shutter.

5. Whitening filters (Pomona box-style) made for the Xarm I and Q channels - these are the same as the whitening for the Y arm.

6. 30m SMA cable made to be used for 2nd delay line.

     6.1 Steve reminded me this morning that we returned one of the fancy spools of cable that was purchased for the delay lines, since it was defective.  We didn't get it replaced because there was debate as to what is the best kind of cable to use.  We need to come to a conclusion, but for now we have a regular RG-405 cable.

7.  Jamie has started work on modifying the beatbox so that we can have 2-arm ALS.  Hopefully that will be done soon-ish, because we're otherwise pretty close to being ready.

  6841   Wed Jun 20 18:43:57 2012 yutaUpdateLSCcalibrated POX error signal

[Jenne, Yuta]

We did the same calibration for POX. It was 3.8e12 counts/m. See elog #6834 for the details of calibration we did.

According to Kiwamu's calibration, actuator response of ITMX is;

A_ITMX  = 4.913e-09 Hz^2*counts/m / freq^2

Plots below are results from our calibration measurement.

LSCxarmTF_usingITMX.pngLSCxarm_HAover1plusG.pngPOXerrorcalibration.png

  6840   Wed Jun 20 18:09:23 2012 yutaUpdateLockingboth arms aligned, ITMX oplev centered

[Jenne, Yuta]

We aligned FPMI. I also centered ITMX oplev because the light was not hitting on QPD.
Alignment procedure we took was;

1. Align Y arm to the Y end green(Y green trans to PSL is now 195 uW with Y end laser measured temperature 34.14 degC).
2. Aligned IR using PZT2 to Yarm(Now, TRY ~ 0.90).
3. Aligned ITMX monitoring AS spots.
4. Aligned X arm so that TRX maximize.
5. Fine adjusted both BS and X arm(Now, TRX ~ 0.82).

Beam spot position on ETMX looks a little too high & left (from ETMXF camera), but we will leave it until ASS scripts is fixed.

FPMIalignment2010620.png

  6839   Wed Jun 20 18:00:10 2012 ranaSummaryPSLsummaries

 Nice PSL summaries from LHO:

https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=3187

  6838   Wed Jun 20 16:37:11 2012 yutaUpdateLockingETMX 1064 trans camera

[Jenne, Yuta]

We made ETMXT camera working.
We connected the camera to video mux, placed 10% pick off mirror in front of TRX PD, lead the beam go to ETMXT camera.
Transmission to the TRY PD was 23.8 uW, but now, it's 21.3 uW (2.3 uW goes to the camera).
So, we changed C1:LSC-TRX_GAIN from -0.00181818 to -0.00203158 (=-0.00181818*23.8/21.3).

There is a channel for power normalization, C1:LSC-TRX_POW_NORM, but is 1 and it looks like we are using this gain for the normalization. Situation of TRY is the same as TRX.

  6837   Wed Jun 20 01:02:20 2012 JamieUpdateGreen LockingRF amp removed from X arm ALS setup

I very badly forgot to log about this in the crush of surfs.

I removed Koji's proto-beatbox RF comparator amp from the X arm ALS setup.  I was investigating hacking it onto one of the discriminator channels in the new beatbox, now that Yuta/Koji's Yuta/Koji's phase tracker is making the coarse beatbox path obsolete.  Upon further reflection we decided to just go ahead and stuff the beatbox board for the X arm, and use the proto-beatbox to test some faster ECL comparators.  This will be done first thing in the morning.

In the meantime the old amp is in my cymac mess on the far left of the electronics bench.

  6836   Wed Jun 20 00:02:16 2012 yutaUpdateGreen Lockingslower scan using phase tracking ALS

For those of you who want to see plots from slower scan.

YarmScan20120619.png

  6835   Wed Jun 20 00:01:04 2012 JenneUpdateLSCcalibrated POY error signal

[Yuta, Jenne]

We have measured the out of loop residual motion of the Yarm while locked with the ALS.  We see ~70pm RMS, as compared to Kiwamu's best of ~24pm RMS.  So we're not yet meeting Kiwamu's best measurement, but we're certainly not in crazy-land.

The Yarm ALS was locked, I took a spectrum of POY11_I_ERR, and used the calibration that we determined earlier this evening.  For reference, I attach a screenshot of our ALS loop filters - we had on all the boosts, and both resonant gain filters (~3Hz and ~16Hz).

A large part of the RMS is coming from the 60Hz power line and the 180Hz harmonic....if we could get rid of these (how were they eliminated from the measurement that Kiwamu used in the paper?? - plotted elog 6780) we would be closer. 

Also, it looks like the hump (in our measurementf ~100Hz, in Kiwamu's ~200Hz) is not quite an order of magnitude higher in amplitude in our measurement vs. Kiwamu's.  We have ~5e-11 m/rtHz, Kiwamu had ~7e-12 m/rtHz.  This increase in noise could be coming from the fact that Yuta and Koji decreased the gain in the Ygreen PDH loop to prevent the PDH box from oscillating. 

While we should still think about why we can't use the same gain that Kiwamu was able to ~6 months ago, we think that we're good enough that we can move on to doing mode scans and residual motion measurements of the Xarm.

 

Attachment 1: LSC_POY_11_I_ERR_calib_19June2012.pdf
LSC_POY_11_I_ERR_calib_19June2012.pdf
Attachment 2: POY_calib_19June2012_FiltBankSettings.png
POY_calib_19June2012_FiltBankSettings.png
  6834   Tue Jun 19 23:36:19 2012 yutaUpdateLSCcalibrated POY error signal

[Jenne, Yuta]

We calibrated POY error signal(C1:LSC-POY11_I_ERR). It was 1.4e12 counts/m.

Modeling of Y arm lock:
  Let's say H is transfer function from Y arm length displacement to POY error signal. This is what we want to measure.
  F is the servo filter (filter module C1:LSC-YARM).
  A is the actuator TF using ITMY. According to Kiwamu's calibration using MICH (see elog #5583),

  A_ITMY  = 4.832e-09 Hz^2*counts/m / freq^2

  We used ITMY to lock Y arm because ITMY is already calibrated.

What we did:
  1. Measured openloop transfer function of Y arm lock using POY error signal using ITMY (G=HFA). We noticed some discrepancy in phase with our model. If we include 1800 usec delay, phase fits well with the measurement. I think this is too big.
LSCyarmTF_usingITMY.png


  2. Measured a transfer function between actuator to POY error signal during lock. This should give us HA/(1+G).
LSCyarm_HAover1plusG.png

  4. Calculated H using measurements above. Assuming there's no frequency dependance in H, we got

  H = 1.4e12 counts/m

POYerrorcalibration.png

 For sanity check; Peak to peak of the POY error signal when crossing the IR resonance is about 800 counts. FWHM is about 1 nm, so our measurement is not so crazy.

  6833   Tue Jun 19 20:26:50 2012 JenneHowToLockingSummer Plan

Jenne and Yuta's Summer Plan

These are the things that we'd like to accomplish, hopefully before Yuta leaves in mid-July

* Yarm mode scan

  ~ Measure residual motion of Yarm cavity when ALS is engaged

* Xarm mode scan

  ~ Align Xarm IR

  ~ Align Xarm green to cavity

  ~ Do mode scan (similar to Yarm)

  ~ Measure residual motion of Xarm cavity when ALS is engaged

* Hold both arms on IR resonance simultaneously (quick proof that we can)

  ~ Modify beatbox so we can use both X and Y at the same time (Jamie will do this Wednesday morning - we've already discussed)

* PRMI + Arms

  ~ Lock the PRMI (which we already know we can do) holding arms off resonance, bring both arms into resonance using ALS

* PRC mode matching - figure out what the deal is

  ~ Look at POP camera with video capture - use software that Eric the Tall wrote with JoeB to measure spot size

* DRMI glitches

  ~ Why can't we keep the DRMI locked stably?

* DRMI + Arms

  ~ Full lock!!

  ~ Make lots of useful diagnostics for aLIGO, measure sensing matricies, etc.

  6832   Mon Jun 18 23:54:31 2012 yutaUpdateGreen Lockingphase tracker for ALS

[Koji, Jenne, Yuta]

Summary:
  We put phase tracker in FINE loop for ALS. We checked it works, and we scanned Y arm by sweeping the phase of the I-Q rotator.
  From the 8 FSR scan using FINE (30 m delay line), we derived that Y arm finesse is 421 +/- 6.

What we did:
  1. We made new phase rotator because current cdsWfsPhase in CDS_PARTS doesn't have phase input. We want to control phase. New phase rotator currently lives in /opt/rtcds/userapps/trunk/isc/c1/models/PHASEROT.mdl. I checked that this works by sweeping the phase input and monitoring the IQ outputs.

  2. We made a phase tracker (/opt/rtcds/userapps/trunk/isc/c1/models/IQLOCK.mdl) and included in c1gcv model. Unit delay is for making a feed back inside the digital system. Currently it is used only for BEATY_FINE (Simulink diagram below). We edited MEDM screens a little accordingly.
newIQLOCK.png


  3. Phase tracking loop has UGF ~ 1.2 kHz, phase margin ~50 deg. They are enough becuase ALS loop has UGF ~ 100 Hz. To control phase tracking loop, use filter module C1:ALS-BEATY_FINE_PHASE (with gain 100). Sometimes, phase tracking loop has large offset because of the integrator and freedom of 360*n in the loop. To relief this, use "CLEAR HISTORY."

  4. Locked Y arm using C1:ALS-BEATY_FINE_PHASE_OUT as an error signal. It worked perfectly and UGF was ~ 90 Hz with gain -8 in C1:ALS-YARM filter module.

  5. Swept phase input to the new phase rotator using excitation point in filter module C1:ALS-BEATY_FINE_OFFSET. Below is the result from this scan. As you can see, we are able to scan for more than the linear range of FINE_I_IN1 signal. We need this extra OFFSET module for scanning because BEATY_FINE_I_ERR stays 0 in the phase tracking loop, and also,  error signal for ALS, output of PHASE module, stays 0 in ALS loop.
YarmScan20120618.png

  6. We analyzed the data from 8FSR scan by FINE with phase tracker using analyzemodescan.py (below). We got Y arm finesse to be 421 +/- 6 (error in 1 sigma). I think the error for the finesse measurement improved because we could done more linear sweep using phase tracker.
fine8FSRscan.png


Next things to do:
  - Phase tracker works amazingly. Maybe we don't need COARSE any more.
  - Install it to X arm and do ALS for both arms.
  - From the series of mode scan we did, mode matching to the arm is OK. There must be something wrong in the PRC, not the input beam. Look into PRC mode matching using video capture and measuring beam size.

  6831   Mon Jun 18 23:38:39 2012 JenneUpdateLSCAdded LSC channels to frames

Since the .ini files get overwritten every time a model is compiled now, we need to put all channels we want saved to frames in the DAQ Channels list inside the model.

I added the _ERR channels for all RFPDs (I and Q for each), as well as the _OUT channels for the DCPDs.  I also added the _OUT channels for the DoF servos (ex. C1:LSC-DARM_OUT).  I don't remember off the top of my head what else we used to save from the LSC model, but those all seemed like ones we'll possibly want access to later. 

We need to go through and do this to all the models we use regularly.

Since SUS hasn't been recompiled in a while, all those channels are saved (until such time as someone does a recompile).  Den has gone through and edited the PEM and OAF .ini files by hand each time he recompiles, so we have that data, although we need to put it into the model (which is the new proper way to acquire channels).

  6830   Mon Jun 18 17:28:03 2012 yutaSummaryComputersbugs in CDS_PARTS/simLinkParts/Fcn

Fcn module in CDS_PARTS is used to include a user defined function in a model.
We should be able to use this by entering desired function, but I found some bugs.

BUG1: Fcn doen't work without ";"

If you put ";" after the function, we can compile.

 sin(u[1]);

But if you put without ";", like

 sin(u[1])

you get the following error message when compiling.

controls@c1ioo
~ 0$ rtcds make c1gcv
### building c1gcv...
Cleaning c1gcv...
Done
Parsing the model c1gcv...
Done
Building EPICS sequencers...
Done
Building front-end Linux kernel module c1gcv...
echo >> target/c1gcvepics/README.making_changes
echo 'Built on date' `date` >> target/c1gcvepics/README.making_changes
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild'

make[1]: Entering directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
make -C /lib/modules/2.6.34.1/build SUBDIRS=/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv modules
make[2]: Entering directory `/usr/src/linux-2.6.34.1-cs'
  CC [M]  /opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o
make[2]: Leaving directory `/usr/src/linux-2.6.34.1-cs'
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c: In function 'feCode':
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c:615: error: expected expression before ';' token
make[3]: *** [/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o] Error 1
make[2]: *** [_module_/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv] Error 2
make[1]: *** [default] Error 2
make: *** [c1gcv] Error 1


BUG2: sindeg doesn't work properly

sindeg should work as cosine with input in degrees.
I made a simple model to test this(below).
model_sindegbug.png


Output of the filter module C1:ALS-BEATY_FINE_PHASE goes to "PHASE_in"
sindeg of this goes to C1:ALS-BEATY_FINE_I_ERR
cosdeg of this goes to C1:ALS-BEATY_FINE_Q_ERR

If you sweep the phase input, you should get sin and cos, but you get the following.
cosdeg (C1:ALS-BEATY_FINE_Q_ERR) looks OK, but sindeg (C1:ALS-BEATY_FINE_I_ERR) looks funny. It looks like ~20000 is its period.

dv_sindegbug.png

  6829   Mon Jun 18 16:23:59 2012 JenneUpdateLockingLSC trigger update

The LSC triggers for the individual filter modules in a filter bank now works.  This is handy so that boosts can come on as soon as a cavity is locked, but will turn off when the cavity unlocks.

You choose which filter modules you want to be triggered, and which ones you want to be manually controlled. 

Example:  LSC-YARM    FM4 and FM5 should always be on, but FM2 and FM3 are controlled by the trigger.  You can set the trigger thresholds for the filter modules independently of the main DoF enable trigger thresholds.

  6828   Mon Jun 18 02:31:43 2012 yutaSummaryGreen Lockinganalysis of mode scan data

I analyzed mode scan data from last week.
Mode matching ratio for Y arm is 86.7 +/- 0.3 %. Assuming we can get rid of TEM01/10 by alignment, this can be improved up to ~ 90%.

Peak search, peak fitting and finnesse calculation:
  I made a python script for doing this. It currently lives in /users/yuta/scripts/modescanresults/analyzemodescan.py.
  What it does is as follows

  1. Read mode scan data(coarse5FSRscan.csv, fine1FSRscan.csv). Each column in the data file should be

[time] [some thing like C1:ALS-BEAT(Y|X)_(COARSE|FINE)_(I|Q)_IN1] [C1:LSC-POY11_I_ERR] [C1:LSC-TRY_OUT]

Each separated by comma. Currently, this script uses only TRY, but it reads all anyway

  2. Find peak in TRY data. For the peak search, it splits data in 1 sec and find local maximum. If the local maximum is higher than given threshold, it recognize it as a peak. If two peaks are very close, it uses higher one. This sometimes fails, because mode scan data we have is not so nice.

  3. Fit each peak with Lorentzian function,

TRY = a*b/(4*(t-c)^2+b^2) + d  (a>0, b>0)

  where a/b is a peak height, b is a linewidth (FWHM), c is a peak position in time, and d is a offset.
  I don't like this, but currently, a/b+c is fixed to the maximum value of TRY data used for fitting. This is because sometimes TRY data is so bad and I couldn't get the peak height correctly. Each points of TRY data doesn't have same error because cavity length is fluctuating and relation between cavity length and TRY is not linear. I think I should use some weighting for the fit, but currently, I just use least squares.

  4. Find TEM00 and calculate FSR in "seconds". I just used "seconds" assuming we did a linear sweep. This script recognize TEM00 from the given threshold.

  5. Calculate finesse using FSR and linewidth of the closest TEM00.

  Below are the result plots from this analysis. Calculated finesse looks quite high (~1000). I think this is from non-linearity in the sweep and error in "measured" line width.
coarse5FSRscan.pngfine1FSRscan.png


Higher order modes and RF sidebands:

  Assuming the curvature of ITMY/ETMY are flat/57.5 m, Y arm length is 38.6 m(FSR 3.9 MHz), positions of HOMs and RF sidebands(11/55 MHz) in frequency domain should look like the plot below.
  The script for calculating this currently lives in /users/yuta/scripts/modescanresults/HOMRFSB.py, inspired by Yoichi's script for KAGRA
HOMRFSB.png

Mode-matching ratio:
  By comparing mode scan data and HOM/RF SB positions in a sophisticated way, you can tell which peak is which.
coarse5FSRscanHOMRFSB.png


  From COARSE 5FSR measurement, peak heights are

TEM00 0.884, 0.896, 0.917, 0.905, 0.911
TEM01 0.040, 0.037, 0.051, 0.054, 0.062
TEM02 0.083, 0.078, 0.079, 0.071, 0.078
TEM03 0.018, 0.015, 0.013, 0.015, 0.014

  So the mode-matching ratio is

MMR = 86.2 %, 87.3 %, 86.5 %, 86.6 %, 85.5 %

  From FINE 1FSR measurement, peak heights and mode matching ratio is

TEM00 0.921
TEM01 0.031
TEM02 0.078
TEM03 0.014

MMR = 88.2 %

  Assuming each measurement had same error, mode-matching ratio from these 6 values is

MMR = 86.7 +/- 0.3 %  (error in 1 sigma)

  This can be improved by ~5% by alignment because we still see ~5% of TEM01/10. Study in systematic errors on going.

  6827   Sat Jun 16 19:32:11 2012 yutaUpdateGreen LockingY arm length using 5FSR scan

I know!
But I think there's some error (~ 10% ?) in calibrating the beatbox. In elog #6815, slope near zero crossing point is about 68 counts/MHz, but now, its 60 counts/MHz. Also, zero crossing point in elog #6815 was 47 MHz, but now, its 45 MHz. 5FSR scan was done between these two calibration measurement.

Quote:

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

 

  6826   Sat Jun 16 18:51:44 2012 KojiUpdateGreen LockingY arm length using 5FSR scan

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

ELOG V3.1.3-